OpenStudy (anonymous):

help! which of the following represents the linear equation 2x=y+3 in slope-intercept form?

OpenStudy (anonymous):

2x=y+3 2x-y=3 divide by 3 \[\frac{ 2x }{ 3 }-\frac{ y }{ 3 }=1\] \[\frac{ x }{\frac{ 3 }{ 2 } }+\frac{ y }{-3 }=1\] it is an equation of line in intercept form. \[ It has x intercept=\frac{ 3 }{ 2 },y intercept=-3\]